A 0.400-kg ice puck, moving east with a speed of 5.72 m/s , has a head-on collision with a 0.950-kg puck initially at rest. Assume that the collision is perfectly elastic. A)What is the speed of the 0.400-kg puck after the collision and what is the direction of the velocity of the 0.400-kg puck after the collision? B) What ise the speed of the 0.950-kg puck after the collision and what is the direction of the velocity of the 0.950-kg puck after the collision?
Answer #1

A)

Velocity, v1' = (0.4 - 0.950) x 5.72/(0.4 + 0.95) = 2.33 m/s

Direction : opposite to initial direction (west)

B)

Velocity, v2' = (2 x 0.4) x 5.72/(0.4 + 0.95) = 3.39 m/s

Direction : east

Comment in case any doubt please rate my answer....